mary was given 1/4 of a sum of money. the remaining money was shared between alice and cody in the ratio 4 : 5. what fraction of the sum of money did cody receive? Give your answer in the simplest form
Answer by stanbon(75887)

mary was given 1/4 of a sum of money. the remaining money was shared between alice and cody in the ratio 4 : 5. what fraction of the sum of money did cody receive? Give your answer in the simplest form
----
Let the "sum of money" be "S".
---
Start::
Mary has S/4
Alice and Cody have (3/4)S
------
Equation:
4x + 5x = (3/4)S
-------
9x = (3/4)S
x = (1/12)S
------
Alice has 4x = 4(1/12)S = (1/3)S
Cody has 5x = 5(1/12)S = (5/12)S
---------------------
Cheers,
Stan H.
---------------